# Brownian motion of a single microparticle partially engulfed by a giant lipid vesicle

### v1

@@ Abstract
The study of interactions between biomimetic membranes and micron-sized particles serves as a fundamental cornerstone for understanding various biological processes. Here, we aim at controlling the engulfment of microparticles in absence of any externally applied force. We investigate systems involving particles and lipid vesicles, where interactions are primarily mediated by electrostatics. Our attention is directed towards dissipative phenomena at the micron and nano scales occurring between a vesicle membrane and a particle, highlighting the influence of the particle penetration depth and the impact of substructures on particle friction. Our work is complemented by hydrodynamic simulations considering the impacts of the shape of the membrane wrapping the particle and the water gap separating the lipid bilayer membrane and the particle on the translational particle drag. Lastly, we explore a system of aqueous microdroplets in oil, stabilized by a single lipid layer and compare the friction experienced by a particle in contact with this monolayer to the friction observed for particles wrapped by lipid bilayers. Surprisingly, the particle drag is higher on the lipid monolayer than the bilayer, which can be explained by introducing a drag contribution due to the partial wetting dynamics of the particle at the monolayer interface.
@@ I INTRODUCTION
Interactions between biological cells and their surrounding medium, such as the internalization of foreign bodies through endocytosis, are crucial for the survival and development of living organisms [1, 2]. However, they are also associated with emerging sanitary threats, such as the ingestion and inhalation of microplastics or nanomaterials, which compromise the integrity of cell membranes [3]. Therefore, understanding the mechanisms of interactions between micron-sized particles and lipid membranes is of paramount importance both for fundamental science and public health. Many biological processes are mediated by specific interactions between a ligand and a receptor (e.g., avidin-biotin complexes or antibody-antigen interactions [4]). Other non-specific interactions also occur in biological environments and have been studied using model systems made of micron-sized particles and giant unilamellar vesicles (GUVs). Spanke et al. tuned particle-membrane adhesion force using the depletion effect by addition of PEG in solution [5]. Ewins et al. demonstrated the electrostatic attraction between negative polystyrene particles and positively charged vesicles mainly composed of DOPC and a few percent of DOTAP, a cationic lipid, and showed that the penetration depth of the particle through the membrane increased with the amount of DOTAP [6].
@@
To be internalized by a vesicle, a particle must overcome the energetic barriers of tension and bending elasticity that resist membrane deformation, and depending respectively on its tension $\sigma$ and bending rigidity $\kappa_{b}$ [7]. When the adhesive energy between a particle and a membrane surpasses these contributions, it is spontaneously engulfed, with a penetration depth depending on the relative magnitudes of these energies.
@@
In this paper, inspired by the work of Ewins et al. [6], we investigate the interaction between (negatively charged) micron-sized spherical particles and (positively charged) DOPC vesicles doped with 5 molDOTAP. We show that an optically trapped particle brought into close proximity with a GUV experiences a force pulling it out of the trap and leading to its spontaneous engulfment. This system exhibits a wide range of equilibrium states, from particle adhesion without optically resolvable deformation of the membrane, to vesicle rupture, encompassing various degrees of particle wrapping. We characterize the translational friction experienced by the particle as a function of its penetration depth and compare our experimental data to a theoretical model built for a particle straddling a monolayer or a membrane bilayer. To obtain a better description of our system, we performed CFD numerical simulations considering three additional parameters: (i) the membrane segment wrapping the particle for three penetration depths, (ii) the presence of a water gap separating the particle and the lipid bilayer, and (iii) a solid substrate in their vicinity. We highlight that an increase of the water gap significantly decreases the particle drag for relatively high membrane viscosities, but has a relatively small effect for low values of the membrane viscosity. Our work also reveals that the presence of membrane substructures (e.g. tubes or buds) close to a particle contributes to increased friction. To discuss the origin and the quantitative measurement of the dissipation due to the lipid bilayers, we also measure the particle translational friction in the presence of a single monolayer. For the latter experiments, we encapsulate silica particles within water droplets in an oil (showing a viscosity similar to water) stabilized by a monolayer of lipids of the same nature. In the monolayer system, surprisingly we observe a significant increase of the particle friction, which is higher than that for particles wrapped by bilayers. This difference in particle drag can be explained considering the different geometries for monolayers and bilayers. In monolayers as in bare fluid interfaces, the particle is partially wetted by two fluids and the monolayer is not following the profile of the protruding particle. Hence, our experiments for monolayers can be explained by a particle drag contribution due to the partial wetting of the particle at the interface, where thermally excited fluctuations of the contact line add an extra contribution to the friction.
@@
Figure 1: (A) Combined fluorescence and bright-field microscopy snapshots of a silica particle (radius $a=0.6\,\mathrm{\mu m}$) initially trapped, escaping from the optical potential well of stiffness $\kappa_{x}=4.6\,\mathrm{pN.\mu m^{-1}}$ and spontaneously penetrating inside a GUV (5 molDOTAP). (B) Typical 1D trajectories of particles spontaneously escaping from the optical trap and penetrating inside a vesicle. The particles are initially confined within a potential well and subsequently extracted as they approach the membrane vicinity (distance below the optical resolution limit). During the wrapping process, the particles deform the membrane and reach an equilibrium position, with a fixed contact line (except for the curve III)
@@ II.1 Dynamics of particle engulfment by a lipid bilayer
In order to characterize the wrapping dynamics of the membrane on the particle, bare silica microspheres (radius $a=0.6\,\mathrm{\mu m}$) were mixed with DOPC GUVs containing 5 molDOTAP. An optical trapping system was then used to bring a particle into the vicinity of a membrane. Most of the time, beyond a critical distance, the attractive force exerted by the vesicle exceeds the restoring force $F_{x}=\kappa_{x}x$ of the trap. This causes the particle to escape and subsequently enter the vesicle, reaching an equilibrium position with a fixed contact line. Fig. 1A shows an example of a particle initially optically trapped in a harmonic potential of stiffness $\kappa_{x}=4.6\,\mathrm{pN.\mu m^{-1}}$, spontaneously and completely engulfed by a vesicle. The wrapping dynamics strongly depends on the membrane tension, as illustrated in Fig. 1B, which displays typical trajectories of particles escaping from the optical trap and penetrating inside a GUV. The curve I shows a particle inducing a localized deformation associated with the opening of a neck, without visible global deformation. About 60 ms after intake, the neck formed by the wrapping of the particle stabilizes (rebound [14]), and the particle reaches an equilibrium position where the particle-membrane contact line is pinned. The particle is completely engulfed by the vesicle, as evidenced by the distance before/after engulfment, approximately equal to $2a$. The curve II indicates a more complex global-scale deformation induced by partial engulfment of the particle, followed by stabilization of the contact line. Finally, the curve III illustrates a globally visible optical deformation of the GUV induced by the particle. About 170 ms after intake, the vesicle relaxes back to its original spherical shape. Partial wrapping of the particle by the membrane leads to an increase in vesicle tension up to its lysis tension, causing its rupture (likely through pore formation). Thus, the particle does not reach an equilibrium position within the membrane in this configuration. Although we could not quantitatively measure the membrane tension for each vesicle, qualitative analysis of vesicle deformation during particle engulfment reveals that significant, globally uniform deformations, indicative of high membrane tension ($\sigma>10^{-6}\,\mathrm{N.m^{-1}}$), lead to a shallower penetration depth compared to vesicles experiencing only local deformations optically undetectable (only related to the opening of a neck). In low-tension vesicles ($\sigma<10^{-6}\,\mathrm{N.m^{-1}}$), subject to thermal shape fluctuations detectable by optical microscopy, the opening of a neck does not induce global deformation due to the area reservoir contained within these undulations. For tensionless GUVs, the only resistance to the strong adhesion between the particle and the membrane comes from the bending constraints, which usually results in the complete engulfment of the particle, however for high-tension vesicles, the particle may reach a state of partial engulfment. In this case, the particle-vesicle configuration is determined by minimizing the total energy of the system $E=E_{w}+E_{\sigma}$, with $E_{w}=2\pi adw$ and $E_{\sigma}=\pi d^{2}\sigma$ representing the adhesion and tension energies, respectively [15]. Thus the penetration depth $d/a$ (defined in Fig. 2A) depends solely on the balance between the particle-membrane adhesion energy $w$, and the tension of the vesicle $\sigma$ with $w=-(d/a)\sigma$. To achieve maximum precision in tracking particle dynamics, measurements at 900 frames per second were conducted. However, even at such high acquisition frequency, it is not feasible to precisely track the particle, and the rate at which it escapes the trap can only be considered as a lower limit, with a velocity $U>440\,\mathrm{\mu m.s^{-1}}$ (curve III) and an associated Stokes’ drag force $F_{drag}=6\pi\eta_{w}aU>5\,\mathrm{pN}$ (with $\eta_{w}=10^{-3}\ \mathrm{Pa.s}$ the viscosity of water). As mentioned earlier, a low-tension vesicle, highly fluctuating, has enough reservoir area to accommodate the uptake of one or more particles [16]. When a vesicle assumes a perfectly spherical shape, with no thermally activated membrane undulations, any further increase in its volume is significantly constrained by its stretching elasticity, limiting bilayer expansion to a few percent before rupture [17]. Thus, the engulfment of a particle into a vesicle increases membrane tension (especially in small vesicles), potentially leading to its disruption. A vesicle rupturing after the uptake of a particle indicates that its tension was initially very high, close to its lysis tension. The condition for half engulfment of a particle by the membrane is $w\geq\sigma$, thus, considering a typical lysis tension value of $\sigma\approx 10^{-2}\,\mathrm{N.m^{-1}}$ [18], the adhesion force $F_{w}\sim aw$ can reach magnitudes on the order of several nN, with $w=10\,\mathrm{mJ}.\mathrm{m}^{-2}$, similar to the surface energy of the biotin-avidin interaction [19]. This analysis highlights the strong adhesion at play, resulting in the wrapping of the particle even in the high tension regime. Particles of different chemical compositions were also tested (polystyrene, melamine formaldehyde), demonstrating that the attraction depends on surface charge. Negative silica and polystyrene particles exhibited a strong affinity for the positively charged vesicles, in contrast to positive melamine resin particles, which showed no adhesion.

### v2

# Diffusion of a single colloid on the surface of a giant vesicle and a droplet

@@ Abstract
The study of interactions between biomimetic membranes and micron-sized particles is crucial for understanding various biological processes. Here, we control microparticle spontaneous engulfment by giant lipid vesicles by tuning particle surface charge, exploring regimes from negligible to strong adhesion. We focus our attention on dissipative phenomena at the micron- and nano- scales, occurring when a particle is wrapped by a lipid vesicle bilayer or when the particle diffuses at the lipid-monolayer interface of a droplet. For particles wrapped by membrane bilayers, we highlight the influence of the particle penetration depth and the impact of substructures on particle friction. Our work is complemented by hydrodynamic simulations that take into account the effects of the shape of the membrane wrapping the particle and the water gap separating the lipid bilayer membrane from the particle on translational particle drag. We show, however, that a purely hydrodynamic model is not suitable to describe the friction of a particle diffusing at the interface of an aqueous microdroplet in oil, stabilized by a single lipid layer. In hydrodynamic models, dissipation is solely described by the surface shear viscosity of the interface and the bulk fluid viscosity, but in this partial wetting configuration, an additional source of dissipation is required to account for fluctuations at the contact line. Hence, through experimental and numerical studies, we demonstrate that the dissipation contributions for the two geometries are fundamentally different.
@@ I INTRODUCTION
Interactions between biological cells and their surrounding medium, such as the internalization of foreign bodies through endocytosis, are crucial for the survival and development of living organisms [1, 2]. However, they are also associated with emerging sanitary threats, such as the ingestion and inhalation of microplastics or nanomaterials, which compromise the integrity of cell membranes [3]. Therefore, understanding the mechanisms of interactions between micron-sized particles and lipid membranes is of paramount importance both for fundamental science and public health, yet, few experiments have been reported in the literature, and most of them focus on systems under very specific conditions.
@@
Ultimately, a vesicle or a cell can be described in terms of the energetic barriers associated with its lipid bilayer: tension $\sigma$ and bending rigidity $\kappa_{b}$ resist membrane deformation and must be overcome for particle wrapping [4]. When the adhesive energy between a particle and a membrane surpasses these contributions, it is spontaneously engulfed, with a penetration depth depending on the relative magnitudes of these energies.
Many biological processes are mediated by specific interactions between a ligand and a receptor (e.g., avidin-biotin complexes or antibody-antigen interactions [5]) and can be investigated using synthetic model systems composed of micron-sized particles and giant unilamellar vesicles (GUVs). For instance, van der Wel et al. [6] explored membrane-mediated particle assemblies using avidin-coated particles and vesicles containing biotinylated lipids. Additionally, non-specific interactions, also prevalent in biological environments, have been examined through similar model systems. Spanke et al. studied the dynamics of particle wrapping by floppy vesicles as a function of PEG concentration, a depletion agent used to trigger adhesion (with adhesion energies of a few $\mathrm{\mu J}.\mathrm{m}^{-2}$) [7]. Their study is set in the context of the bending-dominated regime, where, $a<\lambda_{\sigma}$, ($a$ being the particle radius), with $\lambda_{\sigma}=\sqrt{\kappa_{b}/\sigma}$ as the bendocapillary length, capturing the balance between tension and bending rigidity. Ewins et al. demonstrated the electrostatic attraction between negative polystyrene particles and positively charged vesicles mainly composed of DOPC and a few percent of DOTAP, a cationic lipid, and showed that the penetration depth of the particle through the membrane increased with the amount of DOTAP [8]. Other studies have focused on investigating the interactions of polyelectrolyte-coated colloids with both small and giant lipid vesicles, as reported by Fery et al. [9].
Many theoretical and few experimental studies have been conducted to investigate the behavior of colloids wrapped by lipid membranes. In their pioneering work, Dietrich et al. [10] demonstrated the impact of membrane tension on the penetration depth of latex particles, describing the equilibrium state with a partial wetting model based on tension and adhesion energies, but lacked precise control over the adhesion strength and did not quantitatively explore particle diffusion within the membrane. Dimova et al. [11] performed viscosimetry measurements by analyzing the motion of micron-sized particles on GUVs and derived membrane viscosity using a model assuming a spherical particle straddling a flat membrane at its equator, without accounting for local deformations or the presence of a water gap between the particle and the bilayer. Shigyou et al. [12] studied the diffusion of submicron particles partially (wrapping degree unquantified) or fully engulfed by giant vesicles, where wrapping was obtained by applying strong centrifugation forces. Theoretical and numerical studies [13] have also shown that nanoparticle wrapping is strongly influenced by the particle-membrane adhesion potential range, particle shape, and global vesicle shape changes during wrapping. Theories [14] have also been developed to describe the engulfment of nanoparticles by asymmetric membranes characterized by non-zero spontaneous curvature. Phenomena such as adhesion-induced segregation of membrane components have been considered, and engulfment diagrams have been established.
@@
None of these models account for either the geometry of the membrane wrapping the particle or the water gap separating them. Nevertheless, they have often been used to interpret experiments involving particles partially wrapped by lipid membranes. In this paper, we characterize the translational friction experienced by the particle as a function of its penetration depth and compare our experimental data to a theoretical model built for a particle straddling a monolayer or a membrane bilayer. To obtain a better description of our system, we performed CFD numerical simulations considering three additional parameters: (i) the membrane segment wrapping the particle for three penetration depths, (ii) the presence of a water gap separating the particle and the lipid bilayer, and (iii) a solid substrate in their vicinity. We highlight that an increase of the water gap significantly decreases the particle drag for relatively high membrane viscosities, but has a rather small effect for low values of the membrane viscosity. Our work also reveals that the presence of membrane substructures (e.g. tubes or buds) close to a particle contributes to increased drag. To discuss the origin and the quantitative measurement of the dissipation due to the lipid bilayers, we also measure the particle translational friction in the presence of a single monolayer. For the latter experiments, we encapsulate silica particles within water droplets in an oil (showing a viscosity similar to water) stabilized by a monolayer of lipids of the same nature. In the monolayer system, surprisingly we observe a significant increase of the particle friction, which is higher than that for particles wrapped by bilayers. This difference in particle drag can be explained considering the different geometries for monolayers and bilayers. In monolayers as in bare fluid interfaces, the particle is partially wetted by two fluids and the monolayer is not following the profile of the protruding particle. Hence, our experiments for monolayers can be explained by the partial wetting of the particle at the interface, where thermally excited fluctuations of the contact line add an extra friction contribution. Studying dissipative phenomena at the particle-membrane interface requires initiating adhesion. While particle wrapping by vesicles can occur spontaneously through specific or non-specific interactions, it can also be forced, e.g., by centrifugation. Inspired by the work of Ewins et al. [8], we focus on the charge-induced interaction between micron-sized spherical particles with different zeta potentials and (positively charged) DOPC vesicles doped with 5 mol % DOTAP, exploring regimes from no attachment to strong adhesion. Hence, to tune the adhesion we rely solely on the particle surface potential and the membrane tension, therefore avoiding the use of PEG depletants that are known to affect the membrane properties [21]. Specifically, we focus on negative silica particles and show that in this strong adhesion regime, an optically trapped particle brought close to a GUV experiences a force pulling it out of the trap and leading to its spontaneous engulfment. Theoretical studies and simulations [13] have shown that in the bending-dominated regime, and where both spontaneous curvature and particle-membrane interaction potential range are equal to zero, only unbound and fully wrapped states are stable. Our study, considering high vesicle tension, nonzero potential range, and possible spontaneous curvature, explains the existence of intermediate wrapping states. Hence, this system allows a quantitative analysis of particle friction across a wide range of equilibrium states, from particle adhesion without optically resolvable membrane deformation, to full engulfment.
@@
Figure 1: (A) Combined fluorescence and bright-field microscopy snapshots of a silica particle (radius $a=0.6\,\mathrm{\mu m}$) initially trapped, escaping from the optical potential well of stiffness $\kappa_{x}=4.6\,\mathrm{pN.\mu m^{-1}}$ and spontaneously penetrating inside a GUV (5 molDOTAP). (B) Typical 1D trajectories of particles spontaneously escaping from the optical trap and penetrating inside a vesicle. The particles are initially confined within a potential well and subsequently extracted as they approach the membrane vicinity (distance below the optical resolution limit). During the wrapping process, the particles deform the membrane and reach an equilibrium position, with a fixed contact line (except for the curve III).
Figure 2: Combined fluorescence and bright-field microscopy snapshots showing a positively charged melamine formaldehyde particle (radius $a=1.25\,\mathrm{\mu m}$) under optical trapping. The particle is brought into contact with the membrane but does not spontaneously escape the optical trap (1). A force (black arrow) is applied by translating the vesicle to induce engulfment, but the particle is not stably wrapped, and the deformation is temporary (2). When the vesicle is moved away from the particle using optical tweezers, no adhesion or membrane tube keeps the particle attached to the vesicle (3).
@@ II.1 Dynamics of particle engulfment by a lipid bilayer
In order to characterize the wrapping dynamics of the membrane on the particle, bare silica microspheres (radius $a=0.6\,\mathrm{\mu m}$) were mixed with DOPC GUVs containing 5 molDOTAP. An optical trapping system was then used to bring a particle into the vicinity of a membrane. Most of the time, beyond a critical distance, the attractive force exerted by the vesicle exceeds the restoring force $F_{x}=\kappa_{x}x$ of the trap. This causes the particle to escape and subsequently enter the vesicle, reaching an equilibrium position with a fixed contact line. A similar behavior was also observed by Dietrich et al. [10]. Fig. 1A shows an example of a particle initially optically trapped in a harmonic potential of stiffness $\kappa_{x}=4.6\,\mathrm{pN.\mu m^{-1}}$, spontaneously and completely engulfed by a vesicle.
The wrapping dynamics strongly depends on the membrane tension, as illustrated in Fig. 1B, which displays typical trajectories of particles escaping from the optical trap and penetrating inside a GUV. The curve I shows a particle inducing a localized deformation associated with the opening of a neck, without visible global deformation. About 60 ms after intake, the neck formed by the wrapping of the particle stabilizes (rebound [22]), and the particle reaches an equilibrium position where the particle-membrane contact line is pinned. The particle is completely engulfed by the vesicle, as evidenced by the distance before/after engulfment, approximately equal to $2a$. The curve II indicates a more complex global-scale deformation induced by partial engulfment of the particle, followed by stabilization of the contact line. Finally, the curve III illustrates a globally visible optical deformation of the GUV induced by the particle. About 170 ms after intake, the vesicle relaxes back to its original spherical shape. Partial wrapping of the particle by the membrane leads to an increase in vesicle tension up to its lysis tension, causing its rupture (likely through pore formation). Thus, the particle does not reach an equilibrium position within the membrane in this configuration.
Although we could not quantitatively measure the membrane tension for each vesicle, qualitative analysis of vesicle deformation during particle engulfment reveals that significant, globally uniform deformations (i.e. without micrometric propagation of membrane distortion), indicative of high membrane tension ($\sigma>10^{-6}\,\mathrm{N.m^{-1}}$), lead to a shallower penetration depth compared to vesicles experiencing only local deformations optically undetectable (only related to the opening of a neck). In low-tension vesicles ($\sigma<10^{-6}\,\mathrm{N.m^{-1}}$), subject to thermal shape fluctuations detectable by optical microscopy, the opening of a neck does not induce global deformation due to the area reservoir contained within these undulations.
For tensionless vesicles, resistance to the strong adhesion between the particle and the membrane arises solely from bending constraints, usually leading to complete particle engulfment, however for high-tension GUVs, the particle may undergo partial engulfment. In this case, the particle-vesicle configuration is determined by minimizing the total energy of the system $E=E_{w}+E_{\sigma}$, with $E_{w}=2\pi adw$ and $E_{\sigma}=\pi d^{2}\sigma$ representing the adhesion and tension energies, respectively [23]. Thus the penetration depth $d/a$ (defined in Fig. 3A) depends solely on the balance between the particle-membrane adhesion energy $w$, and the tension of the vesicle $\sigma$ with $w=-(d/a)\sigma$.
To monitor particle dynamics, measurements at 900 FPS were conducted. However, even at such high acquisition frequency, precise particle tracking is not feasible, and the rate at which it escapes the trap only constitutes a lower limit, with a velocity $U>440\,\mathrm{\mu m.s^{-1}}$ (curve III) and an associated Stokes’ drag force $F_{drag}=6\pi\eta_{w}aU>5\,\mathrm{pN}$ (with $\eta_{w}=10^{-3}\ \mathrm{Pa.s}$ the viscosity of water).
As mentioned earlier, a low-tension vesicle, highly fluctuating, has enough reservoir area to accommodate the uptake of one or more particles [24]. When a vesicle assumes a perfectly spherical shape, with no thermally activated membrane undulations, any further increase in its volume is significantly constrained by its stretching elasticity, limiting bilayer expansion to a few percent before rupture [25]. Thus, the engulfment of a particle into a vesicle increases membrane tension (especially in small vesicles), potentially leading to its disruption. A vesicle rupturing after the uptake of a particle indicates that its tension was initially very high, close to its lysis tension. The condition for half engulfment of a particle by the membrane is $w\geq\sigma$, thus, considering a typical lysis tension value of $\sigma\approx 10^{-2}\,\mathrm{N.m^{-1}}$ [26], the adhesion force $F_{w}\sim aw$ can reach magnitudes on the order of several nN, with $w=10\,\mathrm{mJ}.\mathrm{m}^{-2}$, similar to the surface energy of the biotin-avidin interaction [27]. This analysis highlights the strong adhesion at play, resulting in the wrapping of the particle even in the high tension regime. This contrasts with other studies, such as [9], where the wrapping of polyelectrolyte-coated beads by zwitterionic vesicles was observed only for highly fluctuating membranes.
The wrapping trajectories reported by Spanke et al. [22] for particles engulfed by tensionless vesicles (in the bending-dominated regime), by depletion effect, are similar to those observed in our low-tension vesicle experiments (curve I), where the particle rebounds before settling at its final depth. However, the wrapping velocities differ significantly: a few $\mathrm{\mu m.s^{-1}}$ in their case compared to at least several hundred $\mathrm{\mu m.s^{-1}}$ in ours. This suggests that the interaction mechanism (depletion vs. electrostatics) and the associated energies do not substantially influence the overall shape of the wrapping trajectories. In contrast, curves II and III exhibit qualitatively distinct behaviors due to high adhesion and vesicle tension.
To assess whether particle attraction is primarily driven by surface charge, we tested different chemical compositions: silica, polystyrene, aminated silica and melamine formaldehyde. For each material, the particles were brought close to vesicles using optical tweezers, and we observed whether the particle was spontaneously pulled out of the trap by the GUV. Highly negative silica and polystyrene particles were almost systematically engulfed spontaneously (often leading to full wrapping), whereas positive melamine resin particles showed no adhesion, even when a force was applied using the trap (see Fig. 2). Aminated silica particles, weakly negatively charged, exhibited intermediate behavior, with spontaneous engulfment occurring less frequently (see Table 1). Adhesion is thus mainly governed by electrostatics and can be controlled by varying particle surface potential.
Table 1: Correlation between spontaneous engulfment by positively charged vesicles ($\zeta_{\mathrm{DOTAP}}=+33$ mV) and particle zeta potential.
Particle material $\zeta$ (mV) Spontaneous engulfment111“Yes” indicates $>90\%$ of particles were spontaneously engulfed, “No” indicates none, and “Occasional” indicates $<40\%$.
Silica -63 Yes
Polystyrene -54 Yes
Aminated silica -12 Occasional
Melamine formaldehyde +25 No